OpenStudy (anonymous):

Please help! Write a formula for this sequence: 8, -1, -10, -19,... Write a formula for this sequence: 1/5, 3/5, 1, 1 2/5 Find the 3rd term in this sequence: tn = 2n/5 Find the fifth term in this sequence: tn = 12 - 2n

OpenStudy (anonymous):

First question: sequence is a + (n-1)d...where a is the first term, d = difference, so 8 + (n-1)(-9) 8 -9n + 9 or, 17-9n

OpenStudy (anonymous):

Second question: also an arithmetic sequence whose difference is 2/5...so 1/5 + (n-1)(2/5) 1/5 + 2/5 n - 2/5 or, (2/5)n - 1/5

OpenStudy (anonymous):

Third question: Third term of 2n/5 first term, let n = 1, you get 2/5 n = 2, you get 4/5 third term, n = 3, you get 6/5 Third term is 6/5

OpenStudy (anonymous):

Now, you can answer the last question, almost identical to the third question.

OpenStudy (anonymous):

Ooo. I see now. I missed a step in the first two. Thank you so much!

OpenStudy (anonymous):

You're welcome.